RC Meeting, Friday, October 15 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. CHL crash yesterday (about 3 pm) - beam restored 8:20 pm. Efficient beam delivery over OWL shift, DAY shift less productive (<4 ABU)
  2. Hall A pass planning on pass change Monday (2 pass) and then again on Friday (4 pass)
  3. Next week there will be a maintenance day (Wednesday on white board). May be the possibility for (long) access to the hall.

Hall C Status

  1. Efficient data taking - still no significant hall-related downtime
  2. So far we have collected about 1650 pions (as of 7:30 am) in this configuration
  3. Spike in Q2 insulating vacuum [1]. During accelerator down, Jerry and Larry (and Nathan) accessed the hall to possibly pump on insulating vacuum. Local gauge showed excellent vacuum, so left without pumping - pump cart is still in hall in shielded location.
  4. HCCAM02 no longer working. Requires lift to replace (next SAD).

Current Run Plan

  1. Continue data taking until early Nov
  2. Planning a Controlled Access for Monday during pass change - allow Heidi to check HB cryo.

Current Issues

  1. Strange issue with HV controls (SHMS 2Y11- powers off when bringing up the group) [2]. Doesn't seem to happen with new HV gui
  2. Replace flow switches for magnets: next opportunistic access (Joe Beaufait) ( PENDING )
  3. The HB has seen some fluctuations in the LHe levels. At the next opportunistic access Heidi would like to investigate the cryo line. ( PENDING )
    • Heidi believes this is related to a vacuum issue and the access is for diagnostic purposes only. Can be escorted during regular office hours (do not call her in at midnight for this)
    • If indeed it proves to be the case the only remedy would be to warm up the magnet and pump the region (which would be a multiple day operation)
    • So far the PID loops have done their job and Heidi (together with the shift crew) are keeping a close eye on the LHe levels
  4. Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
  5. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  6. Still running with no UPS on the SHMS cryo controls - Steve L. will put in PR for batteries and UPS. Wants to inspect existing system at next opportunity. ( PENDING )

Special Requests

Useful Links